# Data-Driven Loss Functions for Inference-Time Optimization in Text-to-Image Generation

### v1

@@ Abstract
We introduce Learn-to-Steer, a novel framework that learns data-driven objectives for test-time optimization rather than handcrafting them. Our key insight is to train a lightweight classifier that decodes spatial relationships from the diffusion model’s cross-attention maps, then deploy this classifier as a learned loss function during inference. Training such classifiers poses a surprising challenge: they can take shortcuts by detecting linguistic traces rather than learning true spatial patterns. We solve this with a dual-inversion strategy that enforces geometric understanding. Our method dramatically improves spatial accuracy: from 0.20 to 0.61 on FLUX.1-dev and from 0.07 to 0.54 on SD2.1 across standard benchmarks. Moreover, our approach generalizes to multiple relations and significantly improves accuracy.
@@ 1 Introduction
Figure 1: Learn-to-Steer learns how spatial relationships are encoded in attention maps to guide generation. a) Correctly renders all four orientations (above/below/left/right)  b, c) Handles complex scenes with multiple spatial relationships of up to five objects and three relations. Prompts are for illustration purposes only, see  Sec. 4.3 for actual prompt structure.
@@
Spatial relations represent one type of text-to-image misalignment within a broader family of failures, including incorrect attribute binding, entity neglect, and object counting errors. In this work, we focus specifically on 2D spatial relationships (e.g., ”above”, ”below”, ”left of”, ”right of”) and propose a novel framework, Learn-to-Steer, that learns loss functions for test-time optimization. The core of our method is to treat the steering problem as a learning problem. In an offline first step, we train a lightweight classifier that learns to decode spatial relationships directly from the cross-attention maps produced by the diffusion model. This classifier then serves as a data-driven loss function during inference, providing gradients to steer the latent code toward faithful representation of spatial relations. By learning rather than hand-crafting these guidance signals, we can leverage the full complexity of how diffusion models encode spatial relationships across their internal representations, without needing to manually specify which patterns matter.
@@
Our approach reveals a surprising challenge we term “relation leakage”: when naively trained, classifiers can achieve high accuracy by detecting linguistic traces of relation words in attention maps rather than actual spatial patterns. This creates a unique paradox: we need to predict “left of” while remaining invariant to textual traces of “left of”–essentially learning to extract the visual manifestation of spatial concepts while ignoring their linguistic fingerprints. We address this through a simple yet effective data augmentation strategy that forces the classifier to become invariant to these linguistic artifacts and focus solely on the geometric patterns.
@@
Overall, using learned losses for steering generation significantly improves spatial relation accuracy across multiple diffusion models—from 0.07 to 0.54 on SD2.1 and from 0.20 to 0.61 on FLUX.1-dev – while requiring no model fine-tuning, and avoiding the degradation often seen with fine-tuning approaches. Moreover, to the best of our knowledge, we are the first to demonstrate support for multiple simultaneous spatial relations, directly from input prompts. As shown in Fig. 1, it handles prompts with three relations among five objects. We quantify multiple relations generation accuracy using an extended evaluation scheme.
@@
Our main contributions are: (1) We propose test-time steering using a learned loss function rather than handcrafted losses. (2) We identify and solve the “relation leakage” problem, enabling effective training on cross-attention maps. (3) We demonstrate significant improvements over handcrafted losses, on standard benchmarks (GenEval and T2I-CompBench) across four different diffusion models without any model fine-tuning. (4) We handle multiple simultaneous spatial relations in a single image, demonstrating generalization beyond training on single relations. (5) We introduce an extended evaluation scheme for quantifying multiple relation generation, which we will make publicly available.
@@ 2 Related Work
The first strategy involves fine-tuning the model’s weights. Some methods fine-tune the generator, for instance by using curriculum-based contrastive learning (EVOGEN [17]), injecting token-ordering information with a curated spatially-aware dataset (COMPASS [50]), re-captioning datasets with spatial phrases (SPRIGHT [6]), or leveraging chain-of-thought with multimodal LLMs (Image-Generation CoT [52]). Others optimize prompt embeddings directly (RRNet [48]). A related approach involves training lighter adapter modules that inject new information, such as bounding boxes (GLIGEN [30]), LLM-generated layouts (Reason-out-Your-Layout [9]), or conditioning maps like Canny edges [5] (ControlNet [51]). In contrast, our work keeps the original pretrained model frozen, preserving its capabilities.
@@
The second major paradigm, and the one our work builds upon, is test-time optimization. Here, the latent code is steered during the denoising process using a guidance loss. However, the dominant approach is to rely on handcrafted loss functions based on heuristics about how attention maps should behave. This includes methods that impose manually designed constraints on self- or cross-attention maps [7, 1, 38, 29, 35, 16], as well as those using explicit layouts as alignment targets [4, 47, 8, 37]. Some methods optimize only the initial noise $z_{T}$ rather than the full trajectory: InitNO [15] uses handcrafted attention-based losses, while ReNO [10] leverages external reward from human preference models. NPNet [53] takes a different approach by training a network to predict initial noises from a dataset of noise-image pairs filtered by human preference models, focusing on overall image quality rather than specific compositional aspects like spatial relations. Our work presents a fundamental departure: instead of hand-engineering objectives or relying on external models, we learn the function to optimize directly from the model’s own internal representations, allowing us to capture the often complex, patterns of spatial encoding.
@@
Other strategies leverage powerful, external VLMs or MLLMs to provide feedback, using their scores to refine or rank generated images [45, 27, 28]. In contrast, we use a small, lightweight classifier trained on the generator’s own internal signals, avoiding dependence on large, black-box models.
A concurrent work, Flow-GRPO [31], fine-tunes the diffusion model directly using the GenEval [14] benchmark as a supervision signal. Since this supervision is not differentiable, they take a reinforcement learning approach using GRPO. Flow-GRPO achieves high quality generations in several categories in GenEval. Naturally, this fine-tuning method involves more complex training and requires substantially more computational resources than our method.
Multiple Spatial Relations.
@@
Generating scenes with multiple spatial relations is a long-standing challenge. Recent benchmarks on simple geometric shapes confirm that even state-of-the-art text-to-image models struggle with this task [44]. A common strategy is to generate an intermediate layout to guide the process. Early works used GNNs on scene graphs to produce the intermediate layout, and were often limited to simple synthetic shapes [18, 3, 23, 11]. More recent layout-guided methods use LLMs to interpret free-form text, yet their focus is typically on object counts and attributes rather than multiple explicit spatial relations [43, 25]. While compositional diffusion approaches can handle multiple relations, they have primarily been demonstrated on simple synthetic shapes [32]. In contrast, our method demonstrates this capability directly from a text prompt on realistic images, without requiring an explicit intermediate layout.
@@ 4 Method
Figure 2: Classifier training pipeline. Given a spatially-aligned image, we perform dual image inversion (positive and negative prompts) to prevent relation leakage. During denoising, we extract the relevant attention maps and use them to train our classifier.
@@ 4.1 Learning to Decode Relations from Attention
The Relation Leakage Problem.
Consider the seemingly straightforward task of creating training data. Given an image showing “a dog to the left of a cat”, we could invert it using this descriptive prompt, extract the attention maps for “dog” and “cat”, and label them as “left of”. However, our early experiments revealed a critical flaw: classifiers trained this way achieved high accuracy during training but failed when used to guide generation.
Through careful investigation, we discovered the cause: the classifier was learning to take a shortcut. Rather than analyzing the spatial patterns in the attention maps, it was detecting subtle traces of the relation word (“left”) implicitly encoded within them. This phenomenon, which we term relation leakage, meant the classifier could predict “left of” simply by recognizing linguistic artifacts rather than visual evidence.
To verify this we inverted images using deliberately incorrect prompts – for instance, describing a dog that is actually left of a cat as being above it. Remarkably, the classifier predicted “above” rather than the true relation “left of”, confirming that it was reading the prompted relation from the attention maps rather than the actual spatial configuration.
Learning Invariant Spatial Patterns.
@@
To address relation leakage, we need the classifier to learn representations invariant to the traces of the relation word. This presents a unique challenge: unlike typical invariance learning scenarios where we want to ignore a protected attribute while predicting something else  [13, 49, 2], here the “protected attribute” (the relation word traces) and our prediction target (the spatial relation) are semantically the same. We need to predict “left of” while being invariant to linguistic traces of “left of” in the attention maps–essentially learning to extract the visual manifestation of a concept while ignoring its textual fingerprint.
@@
To achieve this invariance, our key insight is to force the classifier to learn from contradictory linguistic signals through a simple, and effective data augmentation strategy. For each training image, we perform inversion twice. First, with a “positive” prompt that correctly describes the spatial relation: “A dog to the left of a cat.” Second, with a “negative” prompt using a randomly sampled incorrect relation: “A dog above a cat.” Crucially, we label both sets of extracted attention maps with the ground-truth relation—“left of”—based on the actual spatial arrangement in the image.
@@
This dual-inversion strategy solves the leakage problem: since the same spatial configuration now appears with different relation words in the prompts, the classifier cannot rely on linguistic shortcuts. It must learn to rely on geometric patterns rather than textual artifacts, thereby achieving invariance to the relation leakage.
@@
The training process is illustrated in Fig. 2, while the full test-time optimization pipeline is shown in Fig. 3. The detailed architecture of our relation classifier is described in Appendix  A.4.
@@
Training Data: Real and Synthetic Images. Our training dataset combines real images from GQA [21] with synthetically generated images from Image-Generation-CoT[52]. Real images provide naturalistic complexity and diverse spatial configurations but yield noisy attention maps when many objects are present. To complement this, we include synthetic images that contain simpler scenes with fewer objects. These images produce cleaner attention maps that resemble and better align with the images encountered during generation.
@@ 4.2 Test-Time Optimization with Learned Steering
During generation, the classifier serves as a learned loss function. Given a prompt, we parse a single relation triplet for the subjection, relation and object $(s,r,o)$; and at timestep $t$, aggregate the corresponding attention maps to obtain $\mathbf{A}_{s}$ and $\mathbf{A}_{o}$. The classifier predicts $\hat{p}=\mathcal{C}(\mathbf{A}_{s},\mathbf{A}_{o},t)$, where $\hat{p}$ is the predicted probability distribution over relation classes. We compute a cross-entropy loss against the desired relation $r$ from the prompt:
$\mathcal{L}_{t}=-\log\hat{p}[r],$
and update the latent code with its gradient: $z_{t}\leftarrow z_{t}-\alpha\nabla_{z_{t}}\mathcal{L}_{t}$. We optimize the first 50% of the denoising steps. In FLUX.1-schnell, we only optimize the initial noise.
@@ 4.4 Implementation Details
We describe here the main implementation details. Additional details are provided in Appendix A.
Cross-Validation. All architectural choices and hyperparameters were selected through systematic validation on the downstream generation task. We created a validation set using the GenEval framework, carefully ensuring no overlap with test prompts. This validation set contains prompts with similar structure but different object combinations and spatial relations.
Data. For Flux-based models, we use 20K real images from GQA and 4K synthetic images. For each image we extract the attention maps roughly every 5 denoising steps, and including the last (cleanest) step e.g. $\{5,10,15,20,25,49\}$ for multi-step models, and in FLUX.1-schnell we use $\{1,3\}$. For SD-based models, we use a smaller set of training data, because the input dimension is lower ($7\times$ less maps in SD) – with a balanced mixture of 4K real and 4K synthetic images. The GQA dataset undergoes preprocessing to reduce ambiguity: we filter out images with multiple instances of the same object category and remove hypernym relations using WordNet [36]. We use inversion techniques to extract the attention maps of the training data for the denoising steps. For Flux-based models, we use RF-Inversion [40] and for SD-based we use Fixed-Point Inversion [42], using a guidance-scale of 0 for both inversion methods.
Relation Extraction. Relation triplets are extracted from the prompt using SpaCy [19] a natural language parsing library, with some hard-coded exceptions, like ”fire hydrant” or ”hot dog”.
@@ 5 Experiments
We compare Learn-to-Steer with strong baselines through quantitative and qualitative evaluation using two benchmark datasets and across four text-to-image diffusion models Flux.1-schnell [26], Flux.1-dev [26], SD2.1 [39] and SD1.4 [39]. Our method demonstrates significant improvements in spatial relations representation than baselines, which often suffer from entity neglect, remaining effective even for atypical scenarios like “A cow above a boy”. Finally, we study the properties of Learn-to-Steer through an ablation study.
@@ 5.1 Quantitative Analysis
Table 1: Comparison of spatial relation scores from the GenEval [14] (G. Pos.) and T2I-CompBench [20] (T. Spatial) benchmarks grouped by base diffusion model. All results were computed using the same test set and hardware, except for COMPASS [50], which we report directly from the official paper. †  indicates methods that use handcrafted spatial loss functions.
Baselines: We compare with several methods. (1) COMPASS [50] fine-tunes the model with spatially-aligned data and a text token ordering module. (2) RRNet [48] optimizes prompt embeddings using spatially-aligned reference images. (3) FOR [24] employs a two-stage test-time approach: first optimizing latents $z_{t}$ with handcrafted losses, then refining initial noise $z_{T}$ using VLM feedback. (4) FOR-Spatial [24] uses only the handcrafted spatial loss component from FOR’s first stage for test-time optimization–since FOR only supports SD1.4 and SD2.1, we adapt this component for Flux models by aggregating cross-attention maps per object (details in Sec. A.1). (5) STORM [16] uses handcrafted test-time losses based on optimal transport for spatial relations. (6) InitNO [15] uses handcrafted attention-based loss functions for test-time optimization, focusing on entity neglect and attribute binding failures. We reproduce results for all methods using their shared code, except COMPASS code was not released–we compare with images from their paper in Appendix B.
@@
Quantitative Results: Table 1 reveals how learned objectives compare to handcrafted ones across four different models. On SD2.1, spatial accuracy rises from 0.07 to 0.54 on GenEval and from 0.089 to 0.365 on T2I-CompBench—transforming a model that barely works for spatial tasks into one that succeeds more than half the time. FLUX.1-dev shows similar patterns, improving from 0.20 to 0.61 and 0.177 to 0.392 respectively. The results highlight two key findings. First, learned objectives consistently outperform handcrafted spatial losses: STORM achieves 0.19 on SD2.1 GenEval while our method reaches 0.54, and our approach shows similar advantages over FOR-Spatial across all tested models. Second, test-time optimization is competitive with state-of-the-art fine-tuning: COMPASS achieves 0.60 on FLUX.1-dev GenEval through model retraining, while our approach reaches 0.61 by steering the frozen model. Importantly, this pattern holds across architectures—from UNet-based SD models to MMDiT-based Flux variants. Notably, while COMPASS improves spatial relations, it degrades other capabilities—on SD2.1 color accuracy drops from 0.85 to 0.71 and counting accuracy falls from 0.44 to 0.20 (Appendix B). Our approach avoids these trade-offs by steering only when needed.
Multiple Spatial Relations: To evaluate generalization to multiple spatial relations, never seen during training, we generated prompts ranging from two relations among three objects to more complex cases with 4–5 objects and up to three relations. Objects were sampled from a set of 17 common object categories (inspired by  [12] and  [41]). We sampled 300 such multi-relation prompts and generated one image per prompt. Since current benchmarks do not support automatic verification of multiple relations, we extended the GenEval framework to support open-vocabulary evaluation. See implementation details in Sec. A.3. Table 2 evaluates how well FLUX.1-dev and our method handle prompts with multiple simultaneous relations. The base model struggles almost completely, with accuracies near 0.01-0.06 across all settings. In contrast, our method achieves substantial gains: 0.28 on 3 objects with 2 relations, and 0.12 and 0.10 on harder 4- and 5-object cases with 3 relations. These results highlight two findings. First, test-time optimization allows our approach to generalize beyond single-relation training, handling prompts with up to three relations and five objects. Second, these results suggest that relations are handled with some degree of independence, allowing our test-time compositional reasoning to remain robust as complexity increases. While the 0.28 accuracy for two relations is lower than the ideal $(0.61)^{2}=0.372$ for perfectly independent relations, the performance degrades predictably rather than failing completely. For instance, the 0.12 accuracy for three relations is closely approximated by extrapolating from the two-relation result ($0.28\times 0.28/0.61=0.128$). See Sec. A.3.1 for the list of objects and Fig. 9 for additional qualitative comparisons to the base model.
Objects Relations FLUX.1-dev + Ours
3 2 0.02 0.28
4 2 0.06 0.24
4 3 0.01 0.12
5 3 0.01 0.10
Table 2: Automatic evaluation of multiple spatial relations. All results were computed using 300 prompts and the same seed across methods. Bold indicates the best result.
@@
Runtime: We analyze the computational costs of our approach. Our method involves an offline training phase followed by test-time optimization during inference. Classifier training requires 7 hours on an NVIDIA A100 80GB GPU. During inference, test-time optimization introduces computational overhead that varies by model: FLUX.1-schnell increases from 0.5s to 16.5s per image, FLUX.1-dev from 11s to 6 minutes, SD2.1 from 4.5s to 90s, and SD1.4 from 4.5s to 80s. While this overhead is substantial, it enables significant spatial accuracy improvements without requiring model fine-tuning.

### v2

@@ Abstract
We introduce Learn-to-Steer, a novel framework that learns data-driven objectives for test-time optimization rather than handcrafting them. Our key insight is to train a lightweight classifier that decodes spatial relationships from the diffusion model’s cross-attention maps, then deploy this classifier as a learned loss function during inference. Training such classifiers poses a surprising challenge: they can take shortcuts by detecting linguistic traces in the cross-attention maps, rather than learning true spatial patterns. We solve this by augmenting our training data with samples generated using prompts with incorrect relation words, which encourages the classifier to avoid linguistic shortcuts and learn spatial patterns from the attention maps. Our method dramatically improves spatial accuracy: from 20% to 61% on FLUX.1-dev and from 7% to 54% on SD2.1 across standard benchmarks. It also generalizes to multiple relations with significantly improved accuracy.
Figure 1: Learn-to-Steer learns how spatial relationships are encoded in attention maps to guide generation. a) Correctly renders all four orientations (above/below/left/right)  b, c) Handles complex scenes with multiple spatial relationships of up to five objects and three relations. Prompts are for illustration purpose, see  Sec. 4.3 for actual prompt structure.
@@ 1 Introduction
Spatial relations represent one type of text-to-image misalignment within a broader family of failures, including incorrect attribute binding, entity neglect, and object counting errors. In this work, we focus specifically on 2D spatial relationships (e.g., ”above”, ”below”, ”left of”, ”right of”) and propose a novel framework, Learn-to-Steer, that learns loss functions for test-time optimization. The core of our method is to treat the steering problem as a learning problem. In an offline first step, we train a lightweight classifier that learns to decode spatial relationships directly from the cross-attention maps produced by the diffusion model. This classifier then serves as a data-driven loss function during inference, providing gradients to steer the latent code toward faithful representation of spatial relations. By learning rather than hand-crafting these guidance signals, we can leverage the complexity of how diffusion models encode spatial relationships across their internal representations, without manually specifying which patterns matter.
@@
Our approach reveals a surprising challenge we term “relation leakage”: when naively trained, classifiers can achieve high accuracy by detecting linguistic traces of relation words in attention maps rather than actual spatial patterns. This creates a unique paradox: we need to predict “left of” while ignoring textual traces of “left of”–essentially learning to extract the visual manifestation of spatial concepts while ignoring their linguistic fingerprints. We address this through a simple yet effective data augmentation strategy that encourages the classifier to ignore these linguistic artifacts.
@@
Overall, using learned losses for steering generation significantly improves spatial relation accuracy across multiple diffusion models—from 7% to 54% on SD2.1 and from 20% to 61% on FLUX.1-dev – while requiring no model fine-tuning, and avoiding the degradation often seen with fine-tuning approaches. Moreover, to the best of our knowledge, we are the first to demonstrate support for multiple simultaneous spatial relations, directly from input prompts. As shown in Fig. 1, it can handle prompts with three relations among five objects. We quantify multiple relations generation accuracy using an extended evaluation scheme.
@@
Our main contributions are: (1) We show how the loss function for test-time steering can be learned from data rather than be handcrafted. (2) We identify and solve the “relation leakage” problem, enabling effective training on cross-attention maps. (3) We demonstrate significant improvements over handcrafted losses, on standard benchmarks (GenEval and T2I-CompBench) across four different diffusion models without any model fine-tuning. (4) We handle multiple simultaneous spatial relations in a single image, demonstrating generalization beyond training on single relations. (5) We introduce an extended evaluation scheme for quantifying multiple relation generation.
@@ 2 Related Work
The first strategy involves fine-tuning the model’s weights. Some methods fine-tune the generator, for instance by using curriculum-based contrastive learning (EVOGEN [19]), injecting token-ordering information with a curated spatially-aware dataset (COMPASS [55]), re-captioning datasets with spatial phrases (SPRIGHT [6]), or leveraging chain-of-thought with multimodal LLMs (Image-Generation CoT [57]). Others optimize prompt embeddings directly (RRNet [53]). A related approach involves training lighter adapter modules that inject new information, such as bounding boxes (GLIGEN [32]), LLM-generated layouts (Reason-out-Your-Layout [9]), or conditioning maps like Canny edges [5] (ControlNet [56]). Similarly, layout-guided methods [10, 58, 11] use explicit spatial layouts of objects as additional input signals to spatially constrain the denoising process. In contrast, our work keeps the original pretrained model frozen and operates directly on text without requiring explicit layout inputs.
@@
The second major paradigm our work builds upon, is test-time optimization. Here, the latent code is steered during the denoising process using a guidance loss. However, the dominant approach is to rely on handcrafted loss functions based on heuristics about how attention maps should behave. This includes methods that impose manually designed constraints on self- or cross-attention maps [7, 1, 44, 31, 36, 18], as well as those using explicit layouts as alignment targets [4, 52, 8, 40]. Some methods optimize only the initial noise $z_{T}$ rather than the full trajectory: InitNO [17] uses handcrafted attention-based losses, while ReNO [12] leverages external reward from human preference models. NPNet [59] takes a different approach by training a network to predict initial noises from a dataset of noise-image pairs filtered by human preference models, focusing on overall image quality rather than specific compositional aspects like spatial relations. Our work presents a fundamental departure: instead of hand-engineering objectives or relying on external models, we learn the function to optimize directly from the model’s own internal representations, capturing the often complex, patterns of spatial encoding.
@@
Other strategies leverage powerful, external VLMs or MLLMs to provide feedback, using their scores to refine or rank generated images [50, 29, 30]. In contrast, we use a lightweight classifier trained on the generator’s own internal signals, avoiding dependence on large, black-box models.
@@
Multiple Spatial Relations. Generating scenes with multiple spatial relations is a long-standing challenge. Recent benchmarks on simple geometric shapes confirm that even state-of-the-art text-to-image models struggle with this task [49]. A common strategy is to generate an intermediate layout to guide the process. Early works used GNNs on scene graphs to produce the intermediate layout, and were often limited to simple synthetic shapes [20, 3, 25, 13]. More recent layout-guided methods use LLMs to interpret free-form text, yet their focus is typically on object counts and attributes rather than multiple explicit spatial relations [48, 27]. While compositional diffusion approaches can handle multiple relations, they have primarily been demonstrated on simple synthetic shapes [33]. In contrast, our method demonstrates this capability directly from a text prompt on realistic images, without requiring an explicit intermediate layout.
@@ 4 Method
Figure 2: Classifier training pipeline. Given a spatially-aligned image, we augment the data with both correct and incorrect relation prompts to prevent relation leakage. During denoising, we extract relevant attention maps to train our classifier.
@@ 4.1 Learning to Decode Relations from Attention
The Relation Leakage Problem. Consider training a classifier on attention maps from inverted images. Given an image of “a dog to the left of a cat”, we could invert it using this prompt, extract attention maps, and label them as “left of”. However, such classifiers achieved high training accuracy but failed during generation. We discovered the cause: instead of analyzing spatial patterns, the classifier detected subtle traces of the relation word (“left”) encoded within the maps. This relation leakage meant the classifier recognized linguistic artifacts rather than visual evidence. For example, when we inverted images using incorrect prompts, like describing a dog actually left of a cat as “above” - the classifier predicted “above” rather than the true “left of”, confirming it was reading the prompted relation from the attention maps rather than the actual spatial configuration.
@@
Addressing relation leakage. We aim to make the relation classifier learn representations that do not depend on linguistic cues from the relation word. This presents a unique challenge: unlike typical invariance learning where we ignore a protected attribute while predicting something else  [14, 54, 2], here the “protected attribute” (relation word traces) and our prediction target (spatial relation) are semantically same. We should predict “left” while ignoring linguistic traces of “left” in the attention maps.
@@
To reduce leakage, we introduce an augmentation strategy to make the classifier ignore the linguistic cues coming from the inversion prompt. For each training image, we perform inversion twice: once with a “mismatching” prompt that correctly describes the spatial relation (e.g., “A dog to the left of a cat”), and once with a prompt containing a randomly sampled incorrect relation (e.g., “A dog above a cat”). Crucially, both sets of extracted attention maps are labeled with the ground-truth relation observed in the image (“left of” in this example). This ensures that both prompts generate positive training samples for the classifier. See  Sec. C.7 for potential limitations.
@@
This augmentation strategy solves the leakage problem: since the same spatial configuration now appears with different relation words in the prompts, the classifier cannot rely on linguistic shortcuts. It must learn to rely on spatial patterns rather than textual artifacts, thereby avoiding relation leakage.
@@
The training process is illustrated in Fig. 2, while the full test-time optimization pipeline is shown in Fig. 3. The detailed architecture of our relation classifier is described in Appendix C.9.
@@
Training Data: Real and Synthetic Images. Our training dataset consists of (1) real images from GQA [23] and (2) synthetically generated images from Image-Generation-CoT[57]. The real images provide naturalistic complexity and diverse spatial configurations but yield noisy attention maps when many objects are present. The synthetic images contain simpler scenes with fewer objects, producing cleaner attention maps that resemble and better align with the images encountered during generation.
@@ 4.2 Test-Time Optimization with Learned Steering
During generation, the classifier serves as a learned loss function. Given a prompt, we parse a single relation triplet for subject, relation and object $(s,r,o)$; and at timestep $t$, aggregate corresponding attention maps to obtain $\mathbf{A}_{s}$ and $\mathbf{A}_{o}$. The classifier predicts $\hat{p}=\mathcal{C}(\mathbf{A}_{s},\mathbf{A}_{o},t)$, where $\hat{p}$ is the predicted probability distribution over relation classes. We compute a cross-entropy loss against the desired relation $r$ from the prompt: $\mathcal{L}_{t}=-\log\hat{p}[r]$, and update the latent code with its gradient: $z_{t}\leftarrow z_{t}-\alpha\nabla_{z_{t}}\mathcal{L}_{t}$. We optimize the first 50% of the denoising steps. In FLUX.1-schnell, we only optimize the initial noise.
@@ 4.4 Implementation Details
We describe here the main implementation details. Additional details are provided in Appendix C.
Cross-Validation. All design choices and hyperparameters were selected through systematic validation on the downstream generation task. We made a validation set using the GenEval framework, carefully ensuring no overlap with test prompts. This set contains prompts with similar structure but different object combinations and spatial relations.
Data. For Flux-based models, we use 20K real images from GQA and 4K synthetic images. For each image we extract the attention maps roughly every 5 denoising steps, and including the last (cleanest) step e.g. $\{5,10,15,20,25,49\}$ for multi-step models, and in FLUX.1-schnell we use $\{1,3\}$. For SD-based models, we use a smaller set of training data, because the input dimension is lower ($7\times$ less maps in SD) – with a balanced mixture of 4K real and 4K synthetic images. The GQA dataset is preprocessed to reduce ambiguity: we filter out images with multiple instances of the same object category and remove hypernym relations using WordNet [37]. We use inversion techniques to extract the attention maps of the training data for the denoising steps. For Flux-based models, we use RF-Inversion [46], for SD we use Fixed-Point Inversion [47], with a guidance-scale of 0 for both inversion methods.
@@ 5 Experiments
We compare Learn-to-Steer with strong baselines through quantitative and qualitative evaluation using two benchmark datasets and across four text-to-image diffusion models: Flux.1-schnell [28], Flux.1-dev [28], SD2.1 [45] and SD1.4 [45]. Our method demonstrates significant improvements in spatial relations representation compared to baselines, which often suffer from entity neglect, remaining effective even for atypical scenarios like “A cow above a boy”. We then comprehensively evaluate generalization and scalability through out-of-distribution categories, compositional scaling to multiple relations, dense prompts, personalized models, and diagonal relations. Finally, we study the properties of Learn-to-Steer through an ablation study.
@@ 5.1 Quantitative Analysis
Table 1: Comparison of spatial relation scores from the GenEval (G. Pos.) and T2I-CompBench (T. Spatial) benchmarks grouped by base diffusion model. All results were evaluated using the same test set, except for COMPASS, which is reported from [55]. COMPASS indicates evaluation using their provided code. †  indicates methods that use handcrafted spatial loss functions.
Method FLUX.1-dev SD 2.1 SD 1.4
Vanilla 0.18 0.15 0.07
COMPASS 0.58 0.52 0.40
Learn-to-Steer (Ours) 0.60 0.60 0.47
Table 2: Out-of-distribution (OOD) performance showing GenEval’s (G. Pos.) scores on categories that are out-of-distribution for our model and COMPASS.
Baselines: We compare with several methods. (1) COMPASS [55] fine-tunes the model with spatially-aligned data and a text token ordering module. (2) RRNet [53] optimizes prompt embeddings using spatially-aligned reference images. (3) FOR [26] employs a two-stage test-time approach: first optimizing latents $z_{t}$ with handcrafted losses, then refining initial noise $z_{T}$ using VLM feedback. (4) FOR-Spatial [26] uses only the handcrafted spatial loss component from FOR’s first stage for test-time optimization–since FOR only supports SD1.4 and SD2.1, we adapt this component for Flux models by aggregating cross-attention maps per object (details in Sec. C.1). (5) STORM [18] uses handcrafted test-time losses based on optimal transport for spatial relations. (6) InitNO [17] uses handcrafted attention-based loss functions for test-time optimization, focusing on entity neglect and attribute binding failures. We reproduce results for all methods using their shared code, except COMPASS code was not released–we compare with images from their paper in Appendix A.1, and in A.2, we further compare to layout-guided methods.
@@
In-distribution categories OOD categories
Obj. Rel. Flux-dev Ours COMPASS Flux-dev Ours COMPASS
3 2 0.05 0.36 0.14 0.02 0.28 0.12
4 2 0.04 0.36 0.21 0.04 0.22 0.14
4 3 0.01 0.20 0.07 0.01 0.19 0.05
5 3 0.01 0.17 0.05 0.01 0.09 0.05
Table 3: Multiple relations accuracy across varying object and relation counts for in-distribution and OOD categories.
@@
Quantitative Results: Table 1 reveals how learned objectives compare to handcrafted ones across four different models. On SD2.1, spatial accuracy rises from 0.07 to 0.54 on GenEval and from 0.089 to 0.365 on T2I-CompBench—transforming a model that barely works for spatial tasks into one that succeeds more than half the time. FLUX.1-dev shows similar patterns, improving from 0.20 to 0.61 and 0.177 to 0.392 respectively. The results highlight two key findings. First, learned objectives consistently outperform handcrafted spatial losses: STORM achieves 0.19 on SD2.1 GenEval while our method reaches 0.54, and our approach shows similar advantages over FOR-Spatial across all tested models. Second, test-time optimization is competitive with state-of-the-art fine-tuning: COMPASS achieves 0.60 on FLUX.1-dev GenEval through model retraining, while our approach reaches 0.61 by steering the frozen model. Importantly, this pattern holds across architectures—from UNet-based SD models to MMDiT-based Flux variants. Notably, while COMPASS improves spatial relations, it degrades other capabilities—on SD2.1 color accuracy drops from 0.85 to 0.71 and counting accuracy falls from 0.44 to 0.20 (Appendix A.1). Our approach avoids these trade-offs by steering only when needed.
@@
Runtime: We analyze the computational costs of our approach. Our method involves an offline training phase followed by test-time optimization during inference. Classifier training requires 7 hours on an NVIDIA A100 80GB GPU. During inference, test-time optimization introduces computational overhead that varies by model: FLUX.1-schnell increases from 0.5s to 16.5s per image, FLUX.1-dev from 11s to 6 minutes, SD2.1 from 4.5s to 90s, and SD1.4 from 4.5s to 80s. While this overhead is substantial, it enables significant spatial accuracy improvements without requiring model fine-tuning. Finally, we analyzed runtime-accuracy tradeoffs for SD 2.1. Performing multiple optimization iterations (15 steps) only during the first 10% of denoising steps while reducing to single optimization steps thereafter cuts runtime from 90s to 25s, while accuracy merely drops from 0.65 to 0.6 on the validation set. Early stages prove most critical - extending this regime to the first 20% of steps (35s) restores accuracy to nearly 0.65.